OpenStudy (anonymous):

Solve this equation for x, 0.69=log x. 1) raise each side of the equation as the power of the base of the log 2)simplify by using the property: z=a^logaz

OpenStudy (anonymous):

anyone?? >:\

OpenStudy (anonymous):

Use this logarithm law: \[\huge \log x=c\] Where x and c are real. This logarithm form can be rewritten as: \[\huge 10^c=x\]

OpenStudy (anonymous):

When there is no base number next to the word log, you take it as base 10.

OpenStudy (anonymous):

But now it asks you to raise both sides. so the log base would be 10.

OpenStudy (anonymous):

well @Azteck so far I have 10^logx=10^0.69 is that right?

OpenStudy (anonymous):

Yes that's correct.

OpenStudy (anonymous):

so what do I do now? >.<

OpenStudy (anonymous):

\[\huge z=a^{\log_{a}z}\] Is that what you mean by the second question?

OpenStudy (anonymous):

blaah I am so confused! I don't know why though. I was understanding everything, but now I am all no where.

OpenStudy (anonymous):

I know I have to use that property to simplify, but I don't know how

OpenStudy (anonymous):

I'm asking you whether the expression given is written as that. I need clarification. This is not confusing as much as trtying to interpret what you wrote in your second question.

OpenStudy (anonymous):

trying*

OpenStudy (anonymous):

ohh yeah that's what it was, sorry!

OpenStudy (anonymous):

Then it's just matching numbers to their corresponding letters... You have this: \[\huge 10^{0.69}=10^{\log x}\] Now you can see that z represents x and a represents 10. \[\huge x=10^{\log_{10}x}\] So then it's just substituting the left hand side of what you came up with int he first question with x. You just replace that whole RHS(Right Hand Side) with x. ANd you get: \[\huge x=10^{0.69}\]

OpenStudy (anonymous):

I meant right hand side sorry. not left hand side.

OpenStudy (anonymous):

Forget what I said about the left hand side and replace that with right hand side.

OpenStudy (anonymous):

in the*

OpenStudy (anonymous):

so, the answer is 4.90?

OpenStudy (anonymous):

\[x=e^{0.69}\]

OpenStudy (anonymous):

so, @Nancy_Lam the definite answer is 1.99?

OpenStudy (anonymous):

You don't find the answer...

OpenStudy (anonymous):

The question says simplify.

OpenStudy (anonymous):

If it said evaluate or solve, then you find the answer.

OpenStudy (anonymous):

let me use the calculator

OpenStudy (anonymous):

umhm. it does say solve for x though...

OpenStudy (anonymous):

Then you should of posted that with your questions. That would of been better for all of us to help you get your answer as quick as possible.

OpenStudy (anonymous):

I did post it, at the beginning...

OpenStudy (anonymous):

And @Nancy_Lam is using natural log. I don't think she has done log to the base e.

OpenStudy (anonymous):

x=1.99372 round = 1.99 you are right

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=0.69%3Dlog+x

OpenStudy (anonymous):

What do you mean @Azteck ??

OpenStudy (anonymous):

If 1.99 is the correct answer your first part will be wrong because you raised both sides by base 10.

OpenStudy (anonymous):

you should of raised both sides by base e if you do that.

OpenStudy (anonymous):

When you get into calculus, everytime you see log, you always assume it is to the base e. But if you haven't gone passed integration etc, then you assume log to be base 10.

OpenStudy (anonymous):

past*

OpenStudy (anonymous):

unless stated otherwise.

OpenStudy (anonymous):

(/.<) ayayayy....

OpenStudy (anonymous):

Have you learnt anything on differentiating e^x? If you haven't done that, then it's unlikely 1.99 is the correct answer.

OpenStudy (anonymous):

I see what you mean... I have to start all over again, I am confused. By the way Thanks ALOT @Azteck & @Nancy_Lam

OpenStudy (anonymous):

No worries.